# On almost commuting unitary matrices

### v1

@@ 2.2. Some general results on orthogonal projections and unitary operators
We also formalize the following known claims: an almost unitary operator is close to a unitary operator, compression of a unitary operator by an almost commuting projection is close to a unitary operator, and that large spectral gaps are stable under such compressions. In what follows, $B(z,R)$ denotes an open ball of radius $R$ about $z\in{\mathbb{C}}$. If $q$ is an orthogonal projection and $a$ is an operator, we will use the following shortened notation for the corresponding compression:
$qa|_{q}:=(qaq)|_{\operatorname{ran}q}.$
@@ 2. $(ii)$
    Suppose that $u\in U(H)$ and $p$ is an orthogonal projection with $\|[p,u]\|\leq\varkappa$. Then there exists a unitary operator $w$ acting on $\operatorname{ran}p$ such that $\|pu|_{p}-w\|\leq 10\varkappa$. Moreover, if $B(z,R)\cap\sigma(u)=\varnothing$, then $B(z,R-10\varkappa)\cap\sigma(w)=\varnothing$.
@@
The first claim follows from the spectral mapping theorem applied to $w^{*}w$. In the second claim, existence follows from applying the first claim to $pu|_{p}$. In order to establish the spectral phenomena, note
$\|(w-z)x\|\geq\|(pup-z)x\|-10\varkappa\|x\|=\|(u-z)x\|-10\varkappa\|x\|\geq(R-10\varkappa)\|x\|,\quad\forall x\in\operatorname{ran}p.$
@@ 2.3. Unitary equivalence between intertwined families of projections
(2.8) $\|r_{j}^{2}-r_{j}\|,\quad\|s_{j}^{2}-s_{j}\|,\quad\|r_{2k-1}r_{2k}\|,\quad\|s_{2k}s_{2k+1}\|,\quad\|r_{j}-s_{j}\|,\quad\|r_{2k-1}+r_{2k}-P_{k}\|,\quad\|s_{2k}+s_{2k+1}-Q_{k}\|$
@@ 3. Isospectral invariant, homotopy, and the winding number
In this section we introduce the isospectral invariant $\operatorname{isospec}(u,v)$, originally defined in [5], state our first main result (the quantitative isospectral homotopy lemma), and use it to show that the isospectral invariant coincides with the winding number invariant.
@@ 3.1. Definition of the isospectral invariant
There exist an absolute constant $C_{1}>0$ such that for any pair of unitary matrices $u,v\in U(n)$ any two intervals $I,J\subset\mathbb{T}$ oriented counterclockwise such that
@@
As a consequence, the condition in equation (3.1) is satisfied for $m$ large enough. Note, however, that all four indicator functions, involved in equation (3.2), commute with each other, and therefore equation (3.2) is well defined even without assuming equation (3.1). Let $e_{1},\ldots,e_{m}$ be the standard basis in ${\mathbb{C}}^{m}$, and let $p_{j}:=\langle e_{j},\cdot\rangle e_{j}$ be the associated projections. If
@@
where $p_{0}:=p_{m}$, and therefore the right hand side of equation (3.2) is always equal to 1-1. We note that, assuming that $m$ is large enough, one can always find a pair of intervals among those in equation (3.4) satisfying equation (3.1). However, the reader can also check that the above calculation can be extended for any pair of intervals oriented counterclockwise, regardless of the norm of the commutator or distance estimates, assuming that the union of these intervals does not contain at least one diagonal entry of $\Omega_{m}$. ∎
@@
For two unitaries $u$ and $v$ with $\operatorname{isospec}(u,v)=0$, the aim of Lemma 1.3 is to construct a piecewise smooth path $v_{t}\colon[0,1]\to U(n)$ satisfying
@@ 3.3. Equivalence between the invariants
We now establish the equivalence of the isospectral invariant with the winding number invariant. This equivalence appears in the literature [11, 5] by identifying each obstruction with the Bott element. However, a more direct proof is possible following the same steps as [5, Section 9], without involving the Bott invariant, as the following quantitative corollary of Lemma 1.3.
@@
Let $u,v\in U(n)$ with $C\|[u,v]\|^{2/5}<1$, where $C$ is the constant from Lemma 1.3. Define $w(u,v)$ to be the winding number of the curve
@@
Remark 3.5.
If $u^{\prime},v^{\prime}]=0$, then, clearly $w(u^{\prime},v^{\prime})=\operatorname{isospec}(u^{\prime},v^{\prime})=0$. As a consequence,
$w(u,v)=\operatorname{isospec}(u,v)=0\quad\text{whenever}\quad\|u-u^{\prime}\|+\|v-v^{\prime}\|\leq\frac{1}{10}.$
We arrive to the well-known fact that the assumption $w(u,v)=0$ is indeed necessary in Theorem 1.2.
@@ 4. Commuting approximants and topological triviality
In this section we prove Theorem 1.4. We will provide a mechanism of how the outcome of Lemma 1.3 allows to reduce the version of Question 1.1 in Theorem 1.2 to another question, which is “topologically equivalent” to Lin’s theorem, where in each step we keep track of quantitative estimates.
4.1. Some preparations
The following proposition combines improved versions of several well-known results, see, for example, see, for example [18, Section 2]. The original proof in [18] relied on the outcome of [17] as a “black box”. In the following proof, which is included mainly for the convenience of the reader, we replace the use of [17] by the stronger result of [21].
@@ 4.1. Some preparations
It is not hard to show that if one replaces $t$ in the last part of Proposition 4.1 by a unitary operator with a spectral gap, the approach would still work. However, the size of the gap will appear in the final estimate. The following lemma quantifies this observation. We note that while it is natural to expect the extra factor $\varepsilon^{-1/2}$, where $\varepsilon$ is the size of the gap, it requires a somewhat delicate estimate of the operator Lipschitz norm of a branch of the argument function, which does not offer a natural rescaling property. This estimate is provided by Lemma A.1 in the Appendix. Modulo that estimate, the proof is very short.
@@
Without loss of generality, one can assume that the spectral gap of $u$ is around 1-1. Let $x:=\frac{1}{2\pi}\arg u=\frac{1}{2\pi}\arg_{\rho}u$ as considered in the Appendix. From Lemma A.1 and (OL6), we have
@@
Lemma 4.2 provides a strategy for proving Theorem 1.2 if one can answer the following question: given two unitary matrices $u,v$ with small commutator, is it possible to “open a gap” in the spectrum of $u$ by a small perturbation, while preserving smallness of the commutator with $v$ in an appropriate (uniform in the dimension and independent of the perturbation size) sense? In view of Remark 3.5, a positive answer to this question must rely on vanishing of the winding number $w(u,v)$. The following result, established (with a short proof) in [24], shows that such gap opening is possible, regardless of the value of $w(u,v)$, if one considers an amplification by $u^{*}$.
@@
Note that, in view of Lemma 4.2, it would be sufficient to transform $u$ and $v$ into a pair of almost commuting unitary matrices where, in addition, one of them has a spectral gap. Following the strategy of retracing the steps of [23], we first consider amplifications of the original pair $(u,v)$ in some space of larger dimension, in which creating a gap would be easier. Afterwards, we carefully descend back into the original space in order to produce a commuting pair. The descent will be performed in two steps, with each step being referred to as “dimension reduction”.
@@
In particular, this allows us to create additional gaps in the spectrum of $u^{\prime}_{\mathrm{amp}}(\varepsilon)$ or split into parts using spectral projections, without further losses in the commutator norm. This idea appears several times in the next subsections.
@@
(4.3) $\|u\oplus u^{\prime}_{\mathrm{path}}({\varepsilon})-u^{\prime}_{\mathrm{amp}}({\varepsilon})\|\leq{\varepsilon}+6{\varepsilon}+C{\varepsilon}^{-1/2}\delta^{1/2},\quad\|[u^{\prime}_{\mathrm{path}}({\varepsilon}),v^{\prime}_{\mathrm{path}}({\varepsilon})]\|=\|[u^{\prime}_{\mathrm{amp}}({\varepsilon}),v^{\prime}_{\mathrm{amp}}({\varepsilon})]\|=0,$
@@
We will use spectral projections of $u\oplus u^{\prime}_{\mathrm{path}}(\varepsilon)$ to shed the extra dimensions of the pair $(u^{\prime}_{\mathrm{amp}}(\varepsilon),v^{\prime}_{\mathrm{amp}}(\varepsilon))$ with the aim of drawing the compressions back into commutation through repeated application of Lemma 4.2. To ease the burden of notation, we will proceed denoting the unitaries above by $u^{\prime}_{\mathrm{amp}},v^{\prime}_{\mathrm{amp}},$ and $u^{\prime}_{\mathrm{path}},v^{\prime}_{\mathrm{path}}$, dropping the dependence on $\varepsilon$.
@@
Assuming that $\gamma<1/100,$ apply Corollary 2.9 to produce a unitary $\sigma\in A_{2d}$ such that
@@
In other words, each of these operators has a large spectral gap, of diameter at least $1/5$. Claim (ii) of Proposition 2.11, applied twice for each of the restrictions (first for the compressions of $w^{\dagger}$ and $v^{\dagger}$, arriving to $g$ and $h$, and then for their compressions (4.7) arriving to to $_{+}, gg_{-}$, $_{+}, hh_{-}$), implies that $g_{+}$ and $g_{-}$ each also inherit spectral gaps of diameter, say, at least $1/5-C\gamma\geq 1/10$, assuming $\gamma<1/1000$.
@@
Then, $g^{\prime}=g_{+}^{\prime}+g_{-}^{\prime}$ and $h^{\prime}=h_{+}^{\prime}+h_{-}^{\prime}$ are unitaries on $\operatorname{ran}(p+q)$, satisfying
@@
Similarly to the previous subsection, the pair $(g^{\dagger},h^{\dagger})$ can now replace the pair $(g^{\prime},h^{\prime})$ for all practical purposes. With this replacement, we gained an additional property that $1_{\{\Re z>1/2\}}(g^{\dagger})\leq p$. We also have the estimates (4.6) with $w^{\dagger},v^{\dagger},s_{-},\gamma$ replaced by $g^{\dagger},h^{\dagger},s_{+},\gamma^{1/2}$, respectively. As a consequence, the compressions
@@
More precisely, we can apply Proposition 2.11 with $\varkappa=\rho=C\gamma^{1/2}\ll 1$ and, in the case of the compressions of $g^{\dagger}$, with the gap diameter $R=1/20$ and center $z=1$ for the compression associated to $s_{+}$ and z=1z=-1 for that corresponding to $(p-s_{+})$.
@@
The commuting unitary matrices $u^{\prime}$, $v^{\prime}$, together with the estimate above, conclude the proof of Theorem 1.4 by noting that

### v2

@@ 2.2. Perturbations of orthogonal projections and unitary operators
We also formalize the following known claims: an almost unitary operator is close to a unitary operator, compression of a unitary operator by an almost commuting projection is close to a unitary operator, and that large spectral gaps are stable under such compressions. In what follows, $B(z,R)$ denotes an open ball of radius $R$ about $z\in{\mathbb{C}}$.
@@ 2. $(ii)$
    Suppose that $u\in U(H)$ and $p$ is an orthogonal projection with $\|[p,u]\|\leq\eta$. Then there exists a unitary operator $w$ acting on $\operatorname{ran}p$ such that $\|pup|_{\operatorname{ran}p}-w\|\leq 10\eta$. Moreover, if $B(z,R)\cap\sigma(u)=\emptyset$, then $B(z,R-10\eta)\cap\sigma(w)=\emptyset$.
@@
The first claim follows from the spectral mapping theorem applied to $w^{*}w$. In the second claim, existence follows from applying the first claim to $pup|_{\operatorname{ran}p}$. In order to establish the spectral phenomena, note
$\|(w-z)x\|\geq\|(pup-z)x\|-10\eta\|x\|=\|(u-z)x\|-10\eta\|x\|\geq(R-10\eta)\|x\|,\quad\forall x\in\operatorname{ran}p.$
@@ 2.3. Unitary equivalence between intertwined families of projections
(2.8) $\|r_{j}^{2}-r_{j}\|,\ \|s_{j}^{2}-s_{j}\|,\ \|r_{2k-1}r_{2k}\|,\ \|s_{2k}s_{2k+1}\|,\ \|r_{j}-s_{j}\|,\ \|r_{2k-1}+r_{2k}-P_{k}\|,\ \|s_{2k}+s_{2k+1}-Q_{k}\|$
@@ 3. Isospectral invariant, homotopy, and the winding number
In this section we discuss the isospectral invariant $\operatorname{isospec}(u,v)$ which was originally defined in [5], state our first main result (the quantitative isospectral homotopy lemma), and use it to show that the isospectral invariant coincides with the winding number invariant.
@@ 3.1. Definition of the isospectral invariant
There exist an absolute constant $C_{1}>0$ such that for any pair of unitary matrices $u,v\in{\mathrm{U}}(n)$ any two intervals $I,J\subset\mathbb{T}$ oriented counterclockwise such that
@@
As a consequence, the condition in equation (3.1) is satisfied for $k$ large enough. Note, however, that all four indicator functions, involved in equation (3.2), commute with each other, and therefore equation (3.2) is well defined even without assuming equation (3.1). Let $e_{1},\ldots,e_{m}$ be the standard basis in ${\mathbb{C}}^{m}$, and let $p_{j}:=\langle e_{j},\cdot\rangle e_{j}$ be the associated projections. If
@@
where $p_{0}:=p_{m}$, and therefore the right hand side of equation (3.2) is always equal to 1-1. We note that, assuming that $k$ is large enough, one can always find a pair of intervals among those in equation (3.4) satisfying equation (3.1). However, the reader can also check that the above calculation can be extended for any pair of intervals oriented counterclockwise, regardless of the norm of the commutator or distance estimates, assuming that the union of these intervals does not contain at least one diagonal entry of $\Omega_{m}$. ∎
@@
For two unitaries $u$ and $v$ with $\operatorname{isospec}(u,v)=0$, the aim of Lemma 1.3 is to construct a piecewise smooth path $v_{t}\colon[0,1]\to{\mathrm{U}}(n)$ satisfying
@@ 3.3. Equivalence between the invariants
We now establish the equivalence of the isospectral invariant with the winding number invariant. This equivalence appears in the literature [11, 5] by identifying each obstruction with the Bott element. However, a more direct proof is possible following the same steps as [5, Section 9], without involving the Bott invariant, as the following quantitative corollary of Lemma 1.3 demonstrates.
@@
Let $u,v\in{\mathrm{U}}(n)$ with $C\|[u,v]\|^{2/5}<1$, where $C$ is the constant from Lemma 1.3. Define $w(u,v)$ to be the winding number of the curve
@@
3.4. Related results: a trace formula and infinite-dimensional amplifications
In [12, Lemma 3.1], it is shown that the winding number invariant can be calculated using the following trace formula
(3.10) $w(u,v)=\frac{1}{2\pi i}\mathrm{Tr}\,\log(vuv^{-1}u^{-1}),$
where $\log\colon\mathbb{C}\setminus(-\infty,0]\to\mathbb{C}$ denotes the principal branch and it is assumed that $\|uv-vu\|<2$. Note that, in the setting of Corollary 3.4 one could also have started with (3.10) as the definition, since it evident that the right-hand-side is a homotopy invariance within the same constraints.
The following results is obtained in [23, Theorem 2.5].
Proposition 3.5.
For every $\varepsilon>0$, there exists $\delta>0$ such that, if $u,v\in\mathrm{U}(n)$ satisfy $\|uv-vu\|<\delta$, then there exist unitary operators $U,V\in\mathrm{U}(\mathbb{C}^{n}\otimes\ell^{2}(\mathbb{N}))$ such that
$\|u\otimes\mathbf{1}-U\|+\|v\otimes\mathbf{1}-V\|<\varepsilon,\quad UV=VU.$
Note that the result does not assume $w(u,v)=0$. In other words, the topological obstruction to obtaining commuting approximants disappears after an infinite amplification. The following lemma demonstrates a similar phenomenon that involves adding an infinite-dimensional identity operator.
Lemma 3.6.
Let $u,v\in{\mathrm{U}}(n)$ with $C\|[u,v]\|^{2/5}<1$, where $C$ is the constant from Lemma 1.3. Then the pair $(u\oplus\mathbf{1}_{\ell^{2}(\mathbb{N})},v\oplus\mathbf{1}_{\ell^{2}(\mathbb{N})})$ is connected to the pair $(\mathbf{1}_{\mathbb{C}^{n}\oplus\ell^{2}(\mathbb{N})},\mathbf{1}_{\mathbb{C}^{n}\oplus\ell^{2}(\mathbb{N})})$ in $\mathrm{U}(\mathbb{C}^{n}\oplus\ell^{2}(\mathbb{N}))$ via a pair of paths $(u_{t},v_{t})$ that satisfies
(3.11) $C\|[u_{t},v_{t}]\|^{2/5}<1,\quad t\in[0,1].$
Proof.
Consider
$U:=(u\oplus u^{*})\oplus(u\oplus u^{*})\oplus\ldots=u\oplus(u^{*}\oplus u)\oplus(u^{*}\oplus u)\ldots\in\mathrm{U}(\mathbb{C}^{n}\oplus\ell^{2}(\mathbb{N}));$
$V:=(v\oplus v^{*})\oplus(v\oplus v^{*})\oplus\ldots=v\oplus(v^{*}\oplus v)\oplus(v^{*}\oplus v)\ldots\in\mathrm{U}(\mathbb{C}^{n}\oplus\ell^{2}(\mathbb{N})).$
Clearly, $\|[U,V]\|=\|[u,v]\|$. Since $w(u,v)=-w(v^{*},u^{*})=-w(u^{*},v^{*})$, the winding number of the pair $(u\oplus u^{*},v\oplus v^{*})$ is zero, and Lemma 1.3 provides a path connecting the pair $(u\oplus u^{*},v\oplus v^{*})$ with $(\mathbf{1}_{{\mathbb{C}}^{2n}},v\oplus v^{*})$. By using any path connecting $v\oplus v^{*}$ with $\mathbf{1}_{{\mathbb{C}}^{2n}}$ in $\mathrm{U}(2n)$, one can see that the pair $(u\oplus u^{*},v\oplus v^{*})$ is connected to $(\mathbf{1}_{{\mathbb{C}}^{2n}},\mathbf{1}_{{\mathbb{C}}^{2n}})$ in $\mathrm{U}(4n)$, with both paths preserving the smallness of the commutators similar to (3.11). Using the first representation for $U$ and $V$, one can transform the pair $(U,V)$ to $(\mathbf{1}_{\mathbb{C}^{n}\oplus\ell^{2}(\mathbb{N})},\mathbf{1}_{\mathbb{C}^{n}\oplus\ell^{2}(\mathbb{N})})$ by moving each copy of the pair $(u\oplus u^{*},v\oplus v^{*})$ along the corresponding copy of the said path. Likewise, the second representation provides the homotopy between $(U,V)$ and the pair $(u\oplus\mathbf{1}_{\ell^{2}(\mathbb{N})},v\oplus\mathbf{1}_{\ell^{2}(\mathbb{N})})$. The proof can be concluded by combining all the above paths, with the estimate (3.11) following from the fact that it holds in each block separately. ∎
We note that the above construction involves a homotopy that affects both $U$ and $V$. While the original homotopy in Lemma 1.3 only transforms $u$, the winding number is invariant under homotopies that affect $u$ and $v$, as long as the commutator remains small. Proposition 3.5 and Lemma 3.6 demonstrate that, in these instances, the winding number obstruction “disappears at infinity”.
@@ 4. Commuting approximants and topological triviality
Figure 2. Noncommutative “Change of Variables,” see Remark 4.2.
In this section we prove Theorem 1.4. By applying the quantitative version of Lin’s theorem we first reduce the problem to the question of opening a uniform gap in the spectrum of one of the unitaries. However, we do not apply this reduction as-is. In order to obtain sharper bounds we apply this reduction to sufficiently large ampliated pairs of matrices, and a proof is furnished by keeping track of quantitative estimates when conducting dimension-reductions of the ampliated pairs along the process.
4.1. Almost commuting matrices with spectral gaps
The following proposition combines improved versions of several well-known results, see, for example [18, Section 2]. The original proof in [18] relied on the outcome of [17] as a “black box”. In the following proof, which is included mainly for the convenience of the reader, we replace the use of [17] by the stronger result of [21].
@@ 4.1. Almost commuting matrices with spectral gaps
Remark 4.2.
It can be helpful to associate each of the cases considered in Proposition 4.1 to a geometric picture of the joint spectra of $s$ and $t$ in the case where their commutator would be exactly zero.
  1. (1)
    In the case $t=t^{*}$ and $s=s^{*}$, assume without loss of generality that $\|s\|\leq 1$ and $\|t\|\leq 1$. Then, the joint spectrum is contained in the square $1,1]\times[-1,1]$.
  2. (2)
    In the case $s=t^{*}$ assume (again, without loss of generality) that $\|s\|=\|t\|\leq 1$. Then, the joint spectrum can be identified with the spectrum of $s$ and is contained in the unit disk.
  3. (3)
    In the case $s=t^{*}$ and $\mathbf{1}\leq|t|\leq 3\cdot\mathbf{1}$, the joint spectrum of is contained in an annulus which is a subset of the (rescaled) disk from the previous case.
  4. (4)
    In the case $t=t^{*}$, $\|t\|\leq 1$, and $s\in\mathrm{U}(n)$, the joint spectrum can be considered as a subset of the cylinder, which is naturally homeomorphic to the annulus from the previous case.
It is not hard to show that if one replaces $t$ in the last part of Proposition 4.1 by a unitary operator with a spectral gap, the approach would still work. However, the size of the gap will appear in the final estimate. The following lemma quantifies this observation. We note that while it is natural to expect the extra factor $\varepsilon^{-1/2}$, where $\varepsilon$ is the size of the gap, it requires a somewhat delicate estimate of the operator Lipschitz norm of a branch of the argument function, which does not offer a natural rescaling property. This estimate is provided by Lemma A.1 in Appendix A. Modulo that estimate, the proof is very short, as illustrated by Figure 2, where the cylinder represents Case (4) in Proposition 4.1.
@@
Without loss of generality, one can assume that the spectral gap of $u$ is around 1-1. Let $x:=\frac{1}{2\pi}\arg u=\frac{1}{2\pi}\arg_{\rho}u$ as considered in Appendix A. From Lemma A.1 and (OL6), we have
@@
Remark 4.4.
In [25], an analogue of Lemma 4.3 was proved for the case where both matrices have spectral gaps. It appears that a modification of the arguments in [25] along the same lines as in Proposition 4.1 ultimately leads to the same conclusion as in Lemma 4.3 (in other words, it is not necessary for the second matrix to have a gap). In particular, [25, Section 2] provides a different construction of a branch of the logarithm with the same properties as the one in Lemma A.1.
A potential strategy for proving Theorem 1.2 would now be to reduce it to Lemma 4.3, which suggest the following question: given two unitary matrices $u,v$ with small commutator, is it possible to “open a gap” in the spectrum of $u$, while preserving smallness of the commutator with $v$ in an appropriate (dimension-independent) sense? Clearly, a meaningful positive answer to this question would provide a path to proving Theorem 1.2, and therefore rely on vanishing of the winding number $w(u,v)$. The following result, established (with a short proof) in [26], shows that such gap opening is possible, regardless of the value of $w(u,v)$, if one is allowed to consider an amplification by $u^{*}$.
Figure 3. When the obstruction vanishes, we decouple red edges.
@@
Remark 4.7.
While the above-mentioned gap opening strategy would work and produce a meaningful version of Theorem 1.2, the actual construction is somewhat more delicate. Lemma 4.3 is first applied to the amplified matrices with $\rho=\delta^{1/3}$ in the beginning of Subsection 4.3. In this case, the pair of unitary operators has a “microscopic gap”, that is, the one whose size goes to 0 as $\delta\to 0$. However, the result of this application allows one to partition the spectrum of one of the matrices in a such a way that further applications involve only “macroscopic gaps” of size that is independent of $\delta$. The case of a macroscopic gap is closer to the kind of gap considered in Proposition 4.1, since it does not require precise control of estimates as the gap size goes to zero, and the commutator losses are the same as in Proposition 4.1, since the factor $\rho^{-1}$ is bounded by a constant.
@@
Note that, in view of Lemma 4.3, it would be sufficient to transform $u$ and $v$ into a pair of almost commuting unitary matrices where, in addition, one of them has a spectral gap. Following the strategy of retracing the steps of [24], we will first consider amplifications of the original pair $(u,v)$ in some space of larger dimension, in which creating a gap would be easier. Afterwards, we carefully descend back into the original space in order to produce a commuting pair. The descent will be performed in two steps, with each step being referred to as “dimension reduction”.
@@
If $q$ is a projection and $a$ is an operator, we will use the following shortened notation for the corresponding compression:
$qa|_{q}:=(qaq)|_{\operatorname{ran}q}.$
@@
In particular, this allows us to create additional gaps in the spectrum of $u^{\prime}_{\mathrm{amp}}(\varepsilon)$ or split into parts using spectral projections, without further losses in the commutator norm. This idea will be used several times in the next subsections.
@@
$p$ $r$ $q$
Figure 4. A “high order” spectral decomposition for $u\oplus u^{\prime}_{\mathrm{path}}(\varepsilon)$ into the unitary $u\in A$ and two semicircular arcs dividing $u^{\prime}_{\mathrm{path}}(\varepsilon)$. After slight modifications (see equations (4.5) and (4.10) below), the spectral projections for the black, red, and blue arcs implement dimension reductions to $u^{\prime}_{\mathrm{amp}}({\varepsilon})$ (with a controlled penalty to commutators).
@@
(4.3) $\|u\oplus u^{\prime}_{\mathrm{path}}({\varepsilon})-u^{\prime}_{\mathrm{amp}}({\varepsilon})\|\leq{\varepsilon}+6{\varepsilon}+C{\varepsilon}^{-1/2}\delta^{1/2},\ \|[u^{\prime}_{\mathrm{path}}({\varepsilon}),v^{\prime}_{\mathrm{path}}({\varepsilon})]\|=\|[u^{\prime}_{\mathrm{amp}}({\varepsilon}),v^{\prime}_{\mathrm{amp}}({\varepsilon})]\|=0,$
@@
We will use spectral projections of $u\oplus u^{\prime}_{\mathrm{path}}(\varepsilon)$ to reduce the extra dimensions of $(u^{\prime}_{\mathrm{amp}}(\varepsilon),v^{\prime}_{\mathrm{amp}}(\varepsilon))$ with the aim of drawing the compressions back into commutation through repeated application of Lemma 4.3. To ease the burden of notation, we will proceed denoting the unitaries above by $u^{\prime}_{\mathrm{amp}},v^{\prime}_{\mathrm{amp}},$ and $u^{\prime}_{\mathrm{path}},v^{\prime}_{\mathrm{path}}$, dropping the dependence on $\varepsilon$.
@@
Assuming that $\gamma<1/100,$ apply Corollary 2.9 to produce a unitary $\sigma\in A^{2d}$ such that
@@
In other words, each of these operators has a large spectral gap, of diameter, say, at least $1/5$. Claim (ii) of Proposition 2.11, applied twice for each of the restrictions (first for the compressions of $w^{\dagger}$ and $v^{\dagger}$, arriving to $g$ and $h$, and then for their compressions (4.7) arriving to to $_{+}, gg_{-}$, $_{+}, hh_{-}$), implies that $g_{+}$ and $g_{-}$ each also inherit spectral gaps of diameter, say, at least $1/5-C\gamma\geq 1/10$, assuming $\gamma<1/1000$.
@@
Then, $g^{\prime}=g_{+}^{\prime}+g_{-}^{\prime}$ and $h^{\prime}=r_{+}^{\prime}+r_{-}^{\prime}$ are unitaries on $\operatorname{ran}(p+q)$, satisfying
@@
Similarly to the previous subsection, the pair $(g^{\dagger},h^{\dagger})$ can now replace the pair $(g^{\prime},h^{\prime})$ for all practical purposes. With this replacement, we gained an additional property that $1_{\{\Re z>1/2\}}(g^{\dagger})\leq p$. We also have the estimates (4.6) with $w^{\dagger},v^{\dagger},s_{-},\gamma$ replaced by $g^{\dagger},h^{\dagger},s_{-},\gamma^{1/2}$, respectively. As a consequence, the compressions
@@
More precisely, we can apply Proposition 2.11 with $\eta=\rho=C\gamma^{1/2}\ll 1$ and, in the case of the compressions of $g^{\dagger}$, with the gap diameter $R=1/20$ and center $z=1$ for the compression associated to $s_{+}$ and z=1z=-1 for that corresponding to $(p-s_{+})$.
@@
The commuting unitary matrices $u^{\prime}$, $v^{\prime}$, together with the estimate above, can now be used to conclude the proof of Theorem 1.4 by noting that